A school principal claims that grade 11 students have a mean grade of 86 with a standard deviation of 4. suppose that the distribution is approximately normal. What is the probability that a random selected grade will be greater than 82 but less that 90?


1
Expert's answer
2021-05-19T16:36:11-0400

Let "X=" grade: "X\sim N(\mu,\sigma^2)"

Given "\mu=86, \sigma=4"

"P(82<X<90)=P(X<90)-P(X\leq82)"

"=P(Z<\dfrac{90-86}{4})-P(Z\leq\dfrac{82-86}{4})"

"=P(Z<1)-P(Z\leq-1)"

"\approx0.84134-0.15866\approx0.6827"